OpenStudy (anonymous):

Find the derivative of f(x) = 7 divided by x at x = 1.

OpenStudy (anonymous):

Do you mean the derivative of f(x)=7/x evaluated at x=1

OpenStudy (anonymous):

Because that would be -7/x^2 at x=1 so -7.

OpenStudy (anonymous):

\[f(x)=\frac{ 7 }{ x }=7x ^{-1},f \prime \left( x \right)=7\left( -1 \right)x ^{-1-1}=-7x ^{-2}=-\frac{ 7 }{ x^2 }\] \[f \prime \left( -1 \right)=\frac{ -7 }{ 1^2 }=-7\]